OpenStudy (anonymous):

What is the value of 2 + (- 2/3)^2 ÷ 1/3 ?

sam (.sam.):

\[\huge \frac{22}{3}\]

sam (.sam.):

\[\frac{2+(-\frac{2}{3})^{2}}{\frac{1}{3}}\] \[=3(2+\frac{4}{9})\] \[=6+\frac{4}{3}\] \[=\frac{22}{3}\]

OpenStudy (anonymous):

0 16 3 -2 These were the answers given.

sam (.sam.):

Is the question \[\huge \frac{2+(-\frac{3}{2})^{2}}{\frac{1}{3}}\]

OpenStudy (anonymous):

Its -2/3 not 3/2

OpenStudy (anonymous):

ok?